OpenStudy (anonymous):

Find the point on the curve y=cos(x) closest to the point (1,1).

OpenStudy (anonymous):

You have two points. A point (x,y) that is on the curve y = cos(x) and a point (1,1) You want (x,y) to be as close as possible to (1,1) In other words. You want the distance between the two points to be a minimum. Minimum! What do we know about minimums? When dy/dx = 0! So first the equation. Distance formula: \[\sqrt{(x-1)^2 + (y-1)^2}\] Now sub in cos x for y \[\sqrt{(x-1)^2 + (\cos x-1)^2}\] Derive this, and set the derivative = to 0. The resulting x value is where the distance is a minimum. Plug in the x into the original equation to find y.

OpenStudy (anonymous):

derivative for that is ridiculous, i have no clue where to go with that thing lol

OpenStudy (anonymous):

The distance of the two points dont necessarily approach 0. The point may not even be on the graph. As in this case it is not, cos(1) = not 1 You have an equation for the Distance D correct? I put that above. You want this distance to be as small as possible, not necessarily 0 (it might in some case). So you want to find the minimum values. You could graph it, and try to visually look at the min (or using a graphing calculator function). But we know, that when the derivative of a function is 0, the function has extrema, in this case a min. So we derive the Distance function, and where the derivative = 0, we have our min,.

OpenStudy (loser66):

@Jeremy0203 which course does this problem come from?

OpenStudy (anonymous):

calc I

OpenStudy (anonymous):

My textbook that covers doing this type of equation is like half a page and doesn't explain in a good way at all how this is done

OpenStudy (anonymous):

Its calc1 yeah. Its optimization. Its abuot problems where you want to find a max or min, with regard to area, distance, volume, prices, things like that. Here this might help http://tutorial.math.lamar.edu/Classes/CalcI/Optimization.aspx

OpenStudy (loser66):

@NobodyOwens that is the max/min for area, not distance!!

OpenStudy (anonymous):

No thats distance too. Here we did the same thing, but we derived the equation for distance. It can be applied to many different things. If you want a distance specific example http://tutorial.math.lamar.edu/Classes/CalcI/MoreOptimization.aspx See Example 3

OpenStudy (anonymous):

It all depends on the problem, but in the end we are deriving to determine max/mins

OpenStudy (loser66):

it seems this problem is over my head!!! http://www.assignmentexpert.com/free-questions/question-on-math-calculus-4905.html

OpenStudy (anonymous):

I got (0.79,0.70)

OpenStudy (anonymous):

I substituted in cos(x) for y in the distance formula. Then I negated the sq rt since minimizing the radical minimizes the whole equation and factored it out and solved for x which was 0.79. Then i put that back into the original equation to solved y this giving me the second point.

OpenStudy (anonymous):

Sound bout right?

OpenStudy (anonymous):

@NobodyOwens

OpenStudy (anonymous):

Yup! And looks right graphicaly too

OpenStudy (anonymous):

awesome. thanks for the help, have a bit better understanding of this now. hopefully i can just get through this calculus class and them fully learn it using resources other than the POS textbook this class uses.

OpenStudy (anonymous):

Allright, all the best luck to you. Ask again if you need help!

OpenStudy (anonymous):

I have another question I do need help with